Do the observed numbers of the genotypes shown below significantly differ from HWE? Do a Chi Square test (1 d.f. X2 = 3.84 at for p < 0.05). Show your table, use 4 significant digits, and clearly state your conclusion.

FF = 36

FS = 38

SS = 26
Answer #1

here proprotion of F allele =(36+36+38)/200=0.55

and that of S allele =1-0.55=0.45

hence expected proportion of FF=0.552 =0.3025

expected proportion of FS =2*0.55*0.45=0.495

expected proportion of SS =0.452 =0.2025

below is chi square goodness of fit test:

           relative observed Expected residual Chi square
category frequency Oi Ei=total*p R2i=(Oi-Ei)/√Ei R2i=(Oi-Ei)2/Ei
FF 0.303 36 30.25 1.05 1.093
FS 0.495 38 49.50 -1.63 2.672
SS 0.203 26 20.25 1.28 1.633
total 1.000 100 100 5.397

as test statistic 5.3974 is higher then critical value 3.84 ; therefore we reject null hypothesis and conclude that observed numbers of the genotypes shown below significantly differ from HWE